OpenStudy (anonymous):

Algebra II help!! what is the solution of the following? -3x-2=-12 9x+6y=-9

OpenStudy (anonymous):

A. (2,1) B. No solutions C. (-2,-1) D. Infinitely many solutions

OpenStudy (mertsj):

Take the first equation and add 2 to both sides. Then divide both sides by -3 Replace x in the second equation with the answer to the first equation and solve for y

OpenStudy (mertsj):

@SamiiBelle What did you get?

OpenStudy (loser66):

the first equation is -3x-2 = -12 or -3x-2y =-12?

OpenStudy (anonymous):

-3x - 2y = -12

OpenStudy (loser66):

no solution, friend

OpenStudy (anonymous):

That's what I thought!! Thanks for verifying!!

OpenStudy (loser66):

ok and you get why , right?

OpenStudy (anonymous):

Sort of. Can you explain it? And help me with one more question?

OpenStudy (loser66):

explain or do another one? or both? if you are willing to study, I am willing to spend time with you. if it's just politeness, please, be honest.

OpenStudy (anonymous):

Both. I want to understand it. I'm very bad at equations.

OpenStudy (loser66):

ok, for this first , divided the second equation by -3 what do you get ( both sides)

OpenStudy (loser66):

hey, it takes that long?

OpenStudy (anonymous):

Sorry, I had to finish my college app. I got -3x+6y=3..is that wrong?

OpenStudy (loser66):

6y /-3 = -2y friend, you miss that term

OpenStudy (loser66):

I re write the first one: -3x -2y =-12 the second one after divided by -3 is -3x-2y = 3 how can? the LHS of 2 equations are the same, but the RHS are NOT the same . Does it make sense? NO, so there is no (x,y) satisfy both them ---> no solutions

OpenStudy (loser66):

got it?

OpenStudy (anonymous):

I do!! Thanks!!

OpenStudy (loser66):

ok, what is next?

OpenStudy (anonymous):

What is the solution of the following system? x-y=11 -x + y=-11

OpenStudy (loser66):

infinite

OpenStudy (anonymous):

Thank you so much! You have saved my life!!